35 g of Na2CO3 is dissolved in 150 mL of water to produce a solution with a total volume of 157.2 mL. If the water is at 33∘C, what would be the freezing point of this solution (in ∘C)
Answer #1

Molality of the solution = 35g of Na2CO3/(molecular wight of Na2CO3*weight of water)=35/(105.99*0.1572kg)= 2.10 molal (m)

One Na2CO3 produces 3 ions. So i = 3

Freezing point depression = Kf*Molality of the solution*i; Kf = 1.86 degC/m * 2.10 m*3 = 11.73 degC

Therefore, freezing point = -11.73 degC